The heights of 220 summer camp attendees are normally distributed with a mean of 62.4 inches and a standard deviation of 3.7 inches. About how many attendees could be expected to be shorter than inches?
A. 5
B. 35
C. 150
D. 215
I'm thinking it's A but not sure. Could someone help?

Answer by stanbon(75887)

The heights of 220 summer camp attendees are normally distributed with a mean of 62.4 inches and a standard deviation of 3.7 inches. About how many attendees could be expected to be shorter than 55 inches?
---------
z(55) = (55-62.4)/3.7 = -2
P(x < 55) = P(z < -2) = normalcdf(-100,-2) = 0.0228
Ans:: 0.0228*220 = 5 when rounded up
--------------
Cheers,
Stan H.
==============
A. 5
B. 35
C. 150
D. 215